OpenStudy (anonymous):

What is the solution? 1/2x^2-x+5=0

OpenStudy (anonymous):

in other words\[\frac{ 1 }{ 2 }x^{2}-x+5=0\] multiple choice \[1\pm i\sqrt{11}\] \[-1\pm i \sqrt{9}\] \[1\pm i \sqrt{9}\] \[-1\pm i \sqrt{11}\]

OpenStudy (anonymous):

since you have answer choices, I would just multiply the roots (ie: (1 + isqrt(11))(1 - isqrt(11)) and see if you get the above equation. hope that helps?